Mela Incorporated obtains fish and then processes them into frozen fillets for distribution to its retail sales department. Direct materials are added at the initiation of the cycle. Conversion costs are incurred evenly throughout the production cycle. Before inspection, some fillets are spoiled due to nondetectible defects. Inspection occurs when units are 50% converted. Spoiled fillets generally constitute 3.5% of the good fillets. Data for April 2008 are as follows:

WIP, beginning inventory 4/1/2008

40,000 fillets

Direct materials (100% complete)

Conversion costs (50% complete)

Started during April

75,000 fillets

Completed and transferred out 4/31/2008

100,000 fillets

WIP, ending inventory 4/31/2008

8,000 fillets

Direct materials (100% complete)

Conversion costs (20% complete)

Costs for April:

WIP, beginning Inventory:

Direct materials

$ 55,000

Conversion costs

40,000

Direct materials added

145,100

Conversion costs added

188,065

What is the number of total spoiled units?

Normal spoilage totals?

What cost is allocated to abnormal spoilage using the weighted-average process-costing method?
Answer #1
Total number of spoiled units:
Units Units
WIP beginning inventory 40000
Add: Stareted during April 75000
To account for 115000
Less:
Completed and transferred out 100000
WIP, ending inventory 8000 108000
Spoiled units 7000
Normal spoilage=Good units completed and transferred out*3.5%=100000*3.5%=3500 units
Cost allocated to Abnormal spoilage:
Abnormal spoilage in units=Total units spoiled-Normal spoilage=7000-3500=3500 units
Direct
Materials
Conversion
costs
Cost of beginning work in process 55000 40000
Cost added during the month 145100 188065
Total costs to account for 200100 228065
Equivalent units of production (Note:1) 115000 108600
Cost per equivalent unit 1.74 2.10
Cost allocated to abnormal spoilage=Units abnormally spoiled*Cost per equivalent unit=3500*(1.74+2.10)=$ 13440
Note:1- Computation of equivalent units of production:
Physical
Units
Direct
Materials
Conversion
costs
Units started and completed 100000 100000 100000
(100000*100%) (100000*100%)
Normal spoilage 3500 3500 3500
(3500*100%) (3500*100%)
Abnormal spoilage 3500 3500 3500
(3500*100%) (3500*100%)
Ending work in process 8000 8000 1600
(8000*100%) (8000*20%)
Equivalent units 115000 108600